# The Landscape of Collapsar Outflows: Structure, Signatures and Origins of Einstein Probe Relativistic Supernova Transients

### v1

@@ 4 Conclusions and Discussion
Einstein Probe is transforming the field of time-domain astrophysics, not only by uncovering new X-ray transients but also by providing precise localization that enables detailed follow-up observations. This capability allows for a more comprehensive understanding of the evolution and diversity of transient phenomena, particularly those linked to collapsars. In this Letter, we build on state-of-the-art numerical simulations to present a coherent picture of the evolution of the various components in collapsars and to connect their observational signatures with stripped-envelope supernovae and the newly discovered Einstein Probe transients.
@@
Observations of stripped-envelope supernovae indicate the presence of a subrelativistic, energetic component that cannot be accounted for by the cocoon alone. We find that the radial structure of the outflow, comprising a mildly relativistic cocoon and subrelativistic ejecta from the black hole accretion disk, matches well the velocity distribution inferred from stripped-envelope supernova observations. The cocoon dominates the outflow at velocities in the range $0.3\lesssim\beta\lesssim 0.95$, and may power the early optical emission through cooling-envelope radiation within the first day. At lower velocities $\beta\lesssim 0.3$, the disk-driven shocked ejecta becomes dominant, likely producing the optical signal at $t\gtrsim 1\,{\rm day}$. At even lower velocities, proto-magnetar outflows may contribute additional cooling emission at later times. Nucleosynthesis in the disk and/or PM outflows enhances the $^{56}{\rm Ni}$ abundance in the outer ejecta layers and gives rise to the bright $^{56}{\rm Ni}$ decay-powered peak.
@@
Based on three Einstein Probe events associated with supernovae but lacking $\gamma$-ray burst detections: EP250108a, EP240414a, and EP250304a, Rastinejad et al. (2025) estimated the event rate of this transient class to be $\sim 10-100\,{\rm Gpc}^{-3}\,{\rm yr}^{-1}$, significantly higher than the $\gamma$-ray burst rate of $\sim 1\,{\rm Gpc}^{-3}\,{\rm yr}^{-1}$ (e.g., Wanderman & Piran, 2010). This higher rate is broadly consistent with a simple cocoon emission scenario, where an axisymmetric jet surrounded by a cocoon in which the bulk of the energy is within an opening angle of $\theta_{c}\approx 0.3$. However, given that $\gamma$-ray burst jets likely wobble, the actual ratio of cocoon-only to jet events is lower. This implies that some events are not powered by cocoon emission, either due to the absence of relativistic outflows (e.g., if the black hole spin is too low) or because the observer is located too far off-axis. In such cases, the optical follow-up may reveal only disk ejecta-driven emission (e.g., in EP250108a), or exclusively supernova emission (as in EP250304a) if, e.g., a magnetized disk (and thus jet, cocoon, and disk ejecta) fails to form.
We conclude that while axisymmetric jet and cocoon models have been extensively studied, collapsars produce a far more complex outflow structure. Numerical simulations highlight the roles of wobbling jets, disk ejecta, and proto-magnetar winds as key components in interpreting stripped-envelope supernovae and Einstein Probe transients. In particular, the comparable energy expected from proto-magnetar outflows and disk ejecta of $\sim 10^{52}\,\,{{\rm erg}}$ is consistent with stripped-envelope supernova observations, raising the question of which component dominates the emission. These insights call for detailed numerical and analytic investigations of the dynamics and emission from each outflow component, from the prompt phase through the late afterglow, and how they may be observationally distinguished from canonical axisymmetric jet-cocoon models.

### v2

@@ 4 Conclusions and Discussion
Einstein Probe is transforming the field of time-domain astrophysics, not only by uncovering new X-ray transients but also by providing precise localization that enables detailed follow-up observations. This capability allows for a more comprehensive understanding of the evolution and diversity of transient phenomena, particularly those linked to collapsars. In this Letter, we build on state-of-the-art numerical simulations to present a coherent picture of the evolution of the various components in collapsars and to connect their observational signatures with stripped-envelope SNe and the newly discovered Einstein Probe transients.
@@
Observations of stripped-envelope SNe indicate the presence of a subrelativistic, energetic component that cannot be accounted for by the cocoon alone. We find that the radial structure of the outflow, comprising a mildly relativistic cocoon and subrelativistic ejecta from the black hole accretion disk, matches well the velocity distribution inferred from stripped-envelope SN observations. The cocoon dominates the outflow at velocities in the range $0.3\lesssim\beta\lesssim 0.95$, and may power the early optical emission through cooling-envelope radiation within the first day. At lower velocities $\beta\lesssim 0.3$, the disk-driven shocked ejecta becomes dominant, likely producing the optical signal at $t\gtrsim 1\,{\rm day}$. At even lower velocities, proto-magnetar outflows may contribute additional cooling emission at later times. Nucleosynthesis in the disk and/or PM outflows enhances the $^{56}{\rm Ni}$ abundance in the outer ejecta layers and gives rise to the bright $^{56}{\rm Ni}$ decay-powered peak.
@@
Based on three Einstein Probe events associated with SNe but lacking $\gamma$-ray burst detections: EP250108a, EP240414a, and EP250304a, Rastinejad et al. (2025) estimated the event rate of this transient class to be $\sim 10-100\,{\rm Gpc}^{-3}\,{\rm yr}^{-1}$, significantly higher than the $\gamma$-ray burst rate of $\sim 1\,{\rm Gpc}^{-3}\,{\rm yr}^{-1}$ (e.g., Wanderman & Piran, 2010). This higher rate is broadly consistent with a simple cocoon emission scenario, where an axisymmetric jet surrounded by a cocoon in which the bulk of the energy is within an opening angle of $\theta_{c}\approx 0.3$. However, given that $\gamma$-ray burst jets likely wobble, the actual ratio of cocoon-only to jet events is lower. This implies that some events are not powered by cocoon emission, either due to the absence of relativistic outflows (e.g., if the black hole spin is too low) or because the observer is located too far off-axis. In such cases, the optical follow-up may reveal only disk ejecta-driven emission (e.g., in EP250108a), or exclusively SN emission (as in EP250304a) if, e.g., a magnetized disk (and thus jet, cocoon, and disk ejecta) fails to form.
We conclude that while axisymmetric jet and cocoon models have been extensively studied, collapsars produce a far more complex outflow structure. Numerical simulations highlight the roles of wobbling jets, disk ejecta, and proto-magnetar winds as key components in interpreting stripped-envelope SNe and Einstein Probe transients. In particular, the comparable energy expected from proto-magnetar outflows and disk ejecta of $\sim 10^{52}\,\,{{\rm erg}}$ is consistent with stripped-envelope SN observations, raising the question of which component dominates the emission. These insights call for detailed numerical and analytic investigations of the dynamics and emission from each outflow component, from the prompt phase through the late afterglow, and how they may be observationally distinguished from canonical axisymmetric jet-cocoon models.